OpenStudy (stamp):

[CALCULUS—OPTIMIZATION] For positive constants A and B, the force between two atoms in a molecule is given by f(r) where r > 0 is the distance between the molecules. What value of r minimizes the force between the atoms?

OpenStudy (stamp):

\[f(r)=-\frac{A}{r^2}+\frac{B}{r^3}\]

OpenStudy (anonymous):

Hello stamp u just need to take the derivative of f(r) and set it equal to zero and solve it for r

OpenStudy (stamp):

That does not work.

OpenStudy (anonymous):

why? can u show me ur work ?

OpenStudy (stamp):

\[f'(r)=\frac{2A}{r^3}-\frac{3B}{r^4}\]

OpenStudy (anonymous):

quite right and what about solving \[f'(r)=0\]?

OpenStudy (stamp):

2Ar - 3B = 0

OpenStudy (anonymous):

and it gives\[r=\frac{3B}{2A}\]

OpenStudy (anonymous):

ok :)